OpenStudy (anonymous):

y=-3x^2-8x -6 solve the parabola

OpenStudy (anonymous):

what do you mean by "solve the parabola"? are you looking for a particular form?

OpenStudy (anonymous):

are you looking for the roots (i.e., y = 0)?

OpenStudy (anonymous):

im looking for the vertex

OpenStudy (anonymous):

okay, so your quadratic is in standard form: y = f(x) = ax^2 + bx +c. the vertex is located at the point x = -b/(2a), y = f(-b/(2a))

OpenStudy (anonymous):

did you find it yet?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did you identify a, b and c?